# System-Theoretic Analysis of Dynamic Generalized Nash Equilibrium Problems – Turnpikes and Dissipativity

### v1

@@ 3.2 A converse turnpike result for GNEPs
3.3 Optimal operation at steady-state GNE
We have established in the previous section under which conditions the open-loop trajectory will spend most of its time at the turnpike, the steady-state GNE. However, this raises the question if the pair $(x_{s},u_{s})$ is also optimal in terms of the entire population performance measure $J(x,u)$. To investigate this, we define optimal steady state operation and suboptimal operation, following [Müller et al., 2015, Def. 2].
Definition 6 (Optimal operation of RHG).
The GNEP (2) is optimally operated at steady-state, if for each $\mathbf{x}\in\mathbb{X}_{0}$ and each game pair $(x,u)\in\mathcal{S}^{\text{\tiny NE}}_{N}(\mathbf{x})$ the following holds $\forall k\in\mathbb{I}_{\geq 0}$:
$\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\ell(x_{k},u_{k})}{N}\geq\ell(x_{s},u_{s}).$ (22)
it is suboptimally operated off steady state if (22) holds in a strict sense.
@@ 3.3 Optimal operation at steady-state GNE
Suppose the set $\mathcal{S}_{\infty}^{GNE}(\mathbf{x})\not$$\emptyset$, then if GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ following Definition 3, it is optimally operated at the steady-state GNE $(x_{s},u_{s})$ and suboptimally operated off steady state.
Proof 3.4.
The proof follows along the lines of [Prop. 6.4][Angeli et al., 2012]. First, we note that $\sup_{x\in\mathbb{X}_{0}}|\Lambda(x)|<\infty$ which allows to state the following equality
@@
From which it follows that
@@
4 Optimality Interpretation of GNEs
In this section we aim to draw a connection between the GNE solution of (2) and the KKT conditions of optimal control problems. In order to do so, we state the following assumptions on the functions in the GNEP (2).
@@ 4 Optimality Interpretation of GNEs
For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J^{v}(\cdot,\cdot,u^{-v})$ and constraint function $g(\cdot,\cdot,u^{-v})$ and the dynamics $f(\cdot,\cdot,u^{-v})$ are continuously differentiable in $x$ and $u^{v}$.
@@
For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J^{v}(\cdot,\cdot,u^{-v})$ is convex and the set $\mathcal{Z}^{v}(\cdot,u^{-v})$ is closed and convex.
The convexity assumption is required to ensure that points which solve the individual agent’s KKT system are not merely stationary points (local minima, saddle points). KKT conditions of GNEs and relevant constraint qualifications have been studied in detail in [Bueno et al., 2019].
@@
$L^{v}_{k}=\begin{bmatrix}1\\ \lambda^{v}_{k+1}\\ \mu_{k}^{v}\\ \nu_{k}^{v}\end{bmatrix}^{\top}\begin{bmatrix}\ell^{v}(x_{k},u_{k},u^{-v}_{k})\\ f(x_{k},u_{k}^{v},u_{k}^{-v})-x_{k+1}\\ g(x_{k},u_{k}^{v},u^{-v}_{k})\\ G^{v}u_{k}^{v}-h^{v}\end{bmatrix}$.
@@
$0\leq\eta_{k}^{v}$ $\quad\perp\quad h^{v}-G^{v}u^{v}_{k}\geq 0.$ (24f)
A well-established result from the GNE literature connects this KKT system to GNEs. The following result appeared as Thm. 4.6 in [Facchinei and Kanzow, 2009].
@@
  2. (b)
    If $(x^{*},u^{*})$ and $(\lambda^{v*},\mu^{v*},\eta^{v*})_{v\in\mathcal{V}}$ solve the KKT system (24) and Ass. 3 holds, then $(x^{*},u^{*})$ is an equilibrium point of the GNEP.
Similar to [Faulwasser and Zanon, 2018], we study when a primal-dual solution to the steady-state GNEP (10) is also a solution for the dynamic counterpart (2). To this end, we consider the Lagrangian for each agent of the steady-state GNE problem in (10) given by
@@
$0\leq\bar{\eta}^{v}$ $\quad\perp\quad h^{v}-G^{v}\bar{u}^{v}\geq 0,$ (25e)
@@ 4.1 The game value function of GNEPs
In the analysis of optimal control problems, the value function plays a fundamental role. It allows to conclusively characterize optimality in Dynamic Programming, it can also be linked to the Maximum Principle. Specifically, in first-order optimality conditions of the Maximum Principle or its simpler Euler-Lagrange variants the gradient of the value function is related to the co-state/adjoint, i.e., to the dual variable linked to the equality constraints imposed by the dynamics. Given that (24) are from a per-agent perspective just first-order (KKT) conditions—though all agents consider the same dynamics and they have to agree on the same sate trajectory—one may wonder what is a GNEP-specific counterpart of the optimal control value function?
To this end, recall the construction of $\ell$ as the sum over all agent-specific costs $\ell^{v}$, cf. (11). Summing $\ell$ from $k=0$ to k=N1k=N-1 leads to the game value function which measures the performance of the agent population at $\mathbf{x}$
$\hskip-4.2679ptV_{N}(\mathbf{x}):=\sum_{k=0}^{N-1}\ell(x_{k}^{*},u^{*}_{k})=\sum_{v\in\mathcal{V}}\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v}).$ (26)
Let $\mathcal{B}_{\varepsilon}(\mathbf{x})\subset\mathbb{R}^{n}$ denote an open ball of radius $\varepsilon$, centered at $\mathbf{x}$.
@@
we have $\mathcal{S}^{GNE}_{N}(\mathbf{x}+\xi)\not$$\emptyset$, i.e., the GNEP (2) admits a solution such that $V_{N}(\mathbf{x})$ is locally differentiable at $\mathbf{x}$.
@@
$\nabla V_{N}(\mathbf{x})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0},$ (27)
@@
Consider the GNEP (24) without inequality constraints and subject to the perturbed initial condition
$x_{0}=\mathbf{x}+\xi,\quad\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})\subset\mathbb{R}^{n}.$
For all $v\in\mathcal{V}$, let $(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))$ denote the solution to the KKT conditions (24) with perturbed initial condition, which for all $\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})$ are optimal due to Theorem 5.
Clearly, it holds that the solution for the unperturbed initial condition is given by
@@
$\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}\left(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)\right)=$ $\phantom{+}\sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial x_{k}}\dfrac{\partial x_{k}}{\partial\xi}$
$+\sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial u^{v}_{k}}\dfrac{\partial u^{v}_{k}}{\partial\xi}$
$+\sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial\lambda^{v}_{k}}\dfrac{\partial\lambda^{v}_{k}}{\partial\xi}$
$+\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi},$
@@
$\left.\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))\right|_{\xi=0}$ $=\left.\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi}\right|_{\xi=0}$
@@
Observe that the definition of $V_{N}(\mathbf{x})$ in (26) implies
$V_{N}(\mathbf{x}+\xi)=\sum_{v\in\mathcal{V}}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi).$
@@
$\dfrac{\mathrm{d}}{\mathrm{d}\xi}V_{N}(\mathbf{x}+\xi)=\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)).$
Evaluation at $\xi=0$ and using (4.1) proves the assertion.
@@ 4.2 Sensitivity analysis of the storage function
The next result translates a crucial observation made in the context of turnpike analysis of optimal control problems to GNEPs. In some sense, the next result provides a storage function counterpart to Theorem 6.
@@
    Any primal-dual solution $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v},\bar{\nu}^{v})$ of (25) constitutes a steady-state solution of (24a24c).
@@
    Let $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v}=0,\bar{\nu}^{v})$ solve (25) such that $(\bar{x},\bar{u})\in\Pi\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ for some $\mathbf{x}\in\mathcal{X}_{0}$. Then, for any storage function $\Lambda$ differentiable at $\bar{x}$, it holds that
@@
with $\ell$ from (11).
Next, we sum (25b) over $v\in\mathcal{V}$ and obtain
@@
Note that Theorem 7 (i) states that (25a25c) is a fixed point for (24a24c) but not for (2) as this would also require the boundary conditions (24d) to hold.
Moreover, the negative sign in (29) stems from the steady-state constraint $f(\bar{x},\bar{u}^{v},\bar{u}^{-v})-\bar{x}=0$ in (10). If this constraint is equivalently written as $\bar{x}-f(\bar{x},\bar{u}^{v},\bar{u}^{-v})=0$, then corresponding Lagrange multiplier flips sign and the minus in (29) can be dropped.
@@
The above result shows two important links between the GNEP (2) and its steady-state counterpart (10): The optimality systems of both problems are closely related which resembles insights from optimal control theory [Trélat and Zuazua, 2015, Zanon and Faulwasser, 2018, Faulwasser and Grüne, 2022]. Moreover, there exists a strong link between the dual variables $\lambda^{v}$ and the storage function $\Lambda$. In the context of optimal control problems similar results have been shown in [Damm et al., 2014, Zanon and Faulwasser, 2018, Faulwasser and Zanon, 2018]. Yet, in the context of GNEPs this link is novel, as the dissipativity-based analysis of GNEPs is still in its infancy.
@@
If the GENP exhibits a turnpike at $(x_{s},u_{s})$ and horizon $N$ is sufficiently large, we arrive at a crucial link between the value function (26) and any differentiable storage function
$\nabla V_{N}(x_{s})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0}\approx\sum_{v\in\mathcal{V}}\lambda_{s}^{v}=-\nabla\Lambda(x_{s})$
or, for all $v\in\mathcal{V}$, $\lambda^{v}_{0}\approx\lambda_{s}^{v}$.

### v3

@@ 3.1 A converse turnpike result for GNEPs
Lastly, the existence of a bounded storage function follows from the observation that the available storage is bounded on $\mathbb{X}_{\infty}(\mathbb{X}_{0})$, cf. Appendix A.
This theorem highlights the close relation of dissipativity and turnpike properties in GNEPs. Theorems 3 and 4 show that under rather mild conditions, strict GNEP dissipativity in the sense of Definition 3 and the turnpike property from Definition 4 are equivalent.
Corollary 5 (Turnpike $\Leftrightarrow$ strict dissipativity).
Consider the GNEP (2) and let Assumptions 13 hold. Suppose that the stage cost $\ell$ is continuous and let $\mathbb{X}_{N}(\mathbb{X}_{0})=\mathbb{X}_{0}$. Then the following two statements are equivalent:
  1. (i)
    The GNEP (2) is strictly dissipative with bounded storage (Definition 3).
  2. (ii)
    The GNEP (2) has the measure turnpike property (Definition 4).
Due to Theorems 3 and 4, the proof is straightforward and thus omitted. The condition $\mathbb{X}_{N}(\mathbb{X}_{0})=\mathbb{X}_{0}$ is imposed to match the set of initial conditions for which Assumption 1 holds with the domain of the available storage. This condition is satisfied if from any point on a GNEP solution trajectory the turnpike $(x_{s},u_{s})$ can be reached in finitely many steps.
3.2 Optimal game operation at steady-state GNEs
We have established in the previous section under which conditions the open-loop trajectory will spend most of its time at the turnpike, the steady-state GNE. However, this raises the question if the pair $(x_{s},u_{s})$ is also the best performing GNE (of potentially very many GNEs) in terms of the entire population performance measure $J_{N}(x,u)$ from (13). To investigate this, we define optimal steady state operation and suboptimal operation, following [47, Def. 2].
Definition 6 (Optimal game operation).
The GNEP (2) is optimally operated at a steady-state GNE, if for each $\mathbf{x}\in\mathbb{X}_{0}$ and each game pair $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ the following holds $\forall k\in\mathbb{I}_{\geq 0}$
$\liminf_{N\to+\infty}\frac{\sum_{k=0}^{N-1}\ell(x_{k},u_{k})}{N}\geq\ell(x_{s},u_{s})$ (28)
for $(x_{s},u_{s})\in\mathcal{S}_{s}^{\text{\tiny GNE}}(\mathbf{x})$. It is suboptimally operated off steady state GNE if (28) holds in a strict sense.
@@ 3.2 Optimal game operation at steady-state GNEs
Suppose that the set $\mathcal{S}_{\infty}^{GNE}(\mathbf{x})\not$$\emptyset$, then if GNEP (2) is strictly dissipative with respect to $(x_{s},u_{s})$ following Definition 3 and with bounded storage, it is optimally operated at the steady-state GNE $(x_{s},u_{s})$ and suboptimally operated off steady state.
Proof 3.5.
The proof follows along the lines of [Prop. 6.4][3]. First, we note that $\sup_{x\in\mathbb{X}_{0}}|\Lambda(x)|<\infty$ giving a bound on the storage function and allowing us to state the following equality
@@
Note that switching to $\liminf$ is necessary as the sequence $\ell(x_{k},u_{k})$ may not be converging. Using superadditivity [^2] Superadditivity: $\liminf\limits_{n\to\infty}(a_{n}+b_{n})\geq\liminf\limits_{n\to\infty}a_{n}+\liminf\limits_{n\to\infty}b_{n}$ of the $\liminf$ and noting that $\ell(x_{s},u_{s})$ is constant we obtain
@@
To the best of our knowledge, Proposition 1 is the first characterization of GNEs from an optimality perspective. It shows that if the group of agents converges to $(x_{s},u_{s})$ (i.e., strict dissipativity holds), then this point is the best possible (hence optimal) steady-state GNE for the population in terms of cost.  [^3] Note that such a characterization exists for potential games [46] that have a direct connection to optimization problems through the potential function. Yet, in (2) we consider a general class of GNEPs.
Figure 2: Overview of implications between strict dissipativity, turnpike, and optimal operation at the steady-state GNE.
4 Value Function and KKT Interpretation
In this section we aim to draw a connection between the GNE solution of (2) and their KKT conditions. In order to do so, we state the following assumptions on the functions in the GNEP (2).
@@ 4 Value Function and KKT Interpretation
For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J_{N}^{v}(\cdot,\cdot,u^{-v})$, the constraint functions $g(\cdot,\cdot,u^{-v}),h^{v}(\cdot),$ and the dynamics $f(\cdot,\cdot,u^{-v})$ are continuously differentiable in $x$ and $u^{v}$.
@@
For every $v\in\mathcal{V}$ and fixed $u^{-v}$, the cost $J^{v}(\cdot,\cdot,u^{-v})$ is convex and the set $\mathcal{Z}_{N}^{v}(\cdot,u^{-v})$ is closed and convex.
The convexity assumption is required to ensure that points which solve the individual agent’s KKT system are not merely stationary points (local minima, saddle points). KKT conditions of GNEs and relevant constraint qualifications have been studied in detail in [9].
@@
$L^{v}_{k}=\begin{bmatrix}1\\ \lambda^{v}_{k+1}\\ \mu_{k}^{v}\\ \nu_{k}^{v}\end{bmatrix}^{\top}\begin{bmatrix}\ell^{v}(x_{k},u_{k},u^{-v}_{k})\\ f(x_{k},u_{k}^{v},u_{k}^{-v})-x_{k+1}\\ g(x_{k},u_{k}^{v},u^{-v}_{k})\\ h^{v}(u_{k}^{v})\end{bmatrix}$.
@@
$0\leq\eta_{k}^{v}$ $\quad\perp\quad-h^{v}(u^{v}_{k})\geq 0.$ (30f)
A well-established result from the GNE literature connects this KKT system to GNEs, cf. [24, Thm. 4.6].
@@
  2. (ii)
    If $(x^{*},u^{*})$ and $(\lambda^{v*},\mu^{v*},\eta^{v*})_{v\in\mathcal{V}}$ solve the KKT system (30) and Ass. 5 holds, then $(x^{*},u^{*})$ is an equilibrium point of the GNEP.
Similar to [29], we study when a primal-dual solution to the steady-state GNEP (11) is also a solution for the dynamic counterpart (2). To this end, we consider the Lagrangian for each agent of the steady-state GNE problem in (11) given by
@@
$0\leq\bar{\eta}^{v}$ $\quad\perp\quad-h^{v}(\bar{u}^{v})\geq 0,$ (31e)
@@ 4.1 The game value function of GNEPs
In the analysis of optimal control problems, the value function plays a fundamental role. It allows to characterize optimality in Dynamic Programming [7], it can also be linked to the Maximum Principle. Specifically, in first-order optimality conditions of the Maximum Principle the gradient of the OCP value function is related to the co-state/adjoint, i.e., to the dual variable linked to the equality constraints imposed by the dynamics. Given that (30) are from a per-agent perspective just first-order (KKT) conditions—though all agents consider the same dynamics and state trajectory—one may wonder is there a GNEP-specific counterpart of the optimal control value function?
To this end, recall the construction of $\ell$ as the sum over all agent-specific costs $\ell^{v}$, cf. (12). Summing $\ell$ from $k=0$ to k=N1k=N-1 leads to the game value function $V_{N}^{*}:\mathbb{X}_{0}\to\mathbb{R}$ which measures the performance of the agent population at $\mathbf{x}$
$\hskip-4.2679ptV^{*}_{N}(\mathbf{x}):=\sum_{k=0}^{N-1}\ell(x_{k}^{*},u^{*}_{k})=\sum_{v\in\mathcal{V}}\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v}).$ (32)
@@
we have $\mathcal{S}^{GNE}_{N}(\mathbf{x}+\xi)\not$$\emptyset$, i.e., the GNEP (2) admits a solution such that $V^{*}_{N}(\mathbf{x})$ is locally differentiable at $\mathbf{x}$.
@@
$\nabla V^{*}_{N}(\mathbf{x})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0},$ (33)
@@
Consider the GNEP (30) without inequality constraints and subject to the perturbed initial condition $x_{0}=\mathbf{x}+\xi,\quad\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})\subset\mathbb{R}^{n}$. For all $v\in\mathcal{V}$, let $(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))$ denote the solution to the KKT conditions (30) with perturbed initial condition, which for all $\xi\in\mathcal{B}_{\varepsilon}(\mathbf{x})$ are dynamic GNEs due to Theorem 6.
The solution for the unperturbed initial condition is
@@
$\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}\left(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)\right)=\\ \sum_{k=0}^{N-1}\dfrac{\partial L_{k}^{v}}{\partial x_{k}}\dfrac{\partial x_{k}}{\partial\xi}+\dfrac{\partial L_{k}^{v}}{\partial u^{v}_{k}}\dfrac{\partial u^{v}_{k}}{\partial\xi}+\dfrac{\partial L_{k}^{v}}{\partial\lambda^{v}_{k}}\dfrac{\partial\lambda^{v}_{k}}{\partial\xi}+\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi},$
@@
$\hskip 0.0pt\left.\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi))\right|_{\xi=0}$ $=\left.\lambda_{0}^{v}(\xi)^{\top}\dfrac{\partial x_{0}}{\partial\xi}\right|_{\xi=0}$
@@
Observe that the definition of $V^{*}_{N}(\mathbf{x})$ in (32) implies
$V^{*}_{N}(\mathbf{x}+\xi)=\sum_{v\in\mathcal{V}}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)).$
@@
$\dfrac{\mathrm{d}}{\mathrm{d}\xi}V^{*}_{N}(\mathbf{x}+\xi)=\sum_{v\in\mathcal{V}}\dfrac{\mathrm{d}}{\mathrm{d}\xi}\mathcal{L}^{v}(x(\xi),u^{v}(\xi),\lambda^{v}(\xi)).$
Setting $\xi=0$ and using (4.1) proves the assertion.
Sensitivity characterization of the value function is an important step towards a more general characterization of value functions in GNEPs with self-interested agents and allows to draw connections to the storage function.
@@ 4.2 Sensitivity analysis of the storage function
The next result translates a crucial observation made in the context of turnpike analysis of OCPs to GNEPs providing a storage function counterpart to Theorem 7.
@@
    Any primal-dual solution $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v},\bar{\nu}^{v})$ of (31) constitutes a steady-state solution of (30a30c).
@@
    Let $(\bar{x},\bar{u},\bar{\lambda}^{v},\bar{\mu}^{v}=0,\bar{\nu}^{v})$ solve (31) and $\bar{x}\in\mathbb{X}_{N}(\mathbb{X}_{0})$ holds. Then, for any storage function $\Lambda$ differentiable at $\bar{x}$, it holds that
@@
with $\ell$ from (12). Now sum (31b) over $v\in\mathcal{V}$ to obtain
@@
Note that Theorem 8 (i) states that (31a31c) is a fixed point for (30a30c) but not for (2) as this would also require the boundary conditions (30d) to hold. Moreover, the negative sign in (35) stems from the steady-state constraint $f(\bar{x},\bar{u}^{v},\bar{u}^{-v})-\bar{x}=0$ in (11). If this constraint is equivalently written as $\bar{x}-f(\bar{x},\bar{u}^{v},\bar{u}^{-v})=0$, then the corresponding Lagrange multiplier flips sign and the minus in (35) can be dropped.
@@
The above result shows two important links between the GNEP (2) and its steady-state counterpart (11): The optimality systems of both problems are closely related which resembles insights from optimal control theory [54, 60, 28]. Moreover, there exists a strong link between the dual variables $\lambda^{v}$ and the storage function $\Lambda$ which has been previously shown for OCPs [18, 60, 29]. Yet, to the best of our knowledge in the context of GNEPs this link is novel.
@@
If the GNEP exhibits a turnpike at $(x_{s},u_{s})$ and the horizon $N$ is sufficiently large, we arrive at a crucial link between the value function (32) and any differentiable storage function
$\nabla V^{*}_{N}(x_{s})=\sum_{v\in\mathcal{V}}\lambda^{v}_{0}\approx\sum_{v\in\mathcal{V}}\lambda_{s}^{v}=-\nabla\Lambda(x_{s})$
or, for all $v\in\mathcal{V}$, $\lambda^{v}_{0}\approx\lambda_{s}^{v}$. The horizon being sufficiently large ensures that the GNEP trajectory converges sufficiently close to $x_{s}$.